Suppose the income for Themba is R40 per month. Themba decides to use all his income on consuming two goods which are fish and pork. The price of fish is R5 and the price of pork is R8.

a)   Clearly illustrate on the graph the combination of the two goods he can afford. Let fish be on the vertical axis and pork on the horizontal axis.

b)   Calculate the slope of the budget line.

 

Question 3

Suppose electricity is a fixed cost and the price and the price of electricity rises.

       I.        What happens to the average total cost curve (ATC)? and illustrate this effect graphically.

      II.        Graphically illustrate what happens on the marginal cost curve.

 

 


1
Expert's answer
2021-06-15T12:54:41-0400

Budget line is the locus of various combination of two goods which a consumer can purchase with given income and given price. 

Given : 

Income = R40

Price of Fish = R5

Price of Pork = R8

A budget line is represented as XPx+YPy = M

where M is the income,

Px is the price of good x

Py is the price of good y

X is the quantity of good X

Y is the quantity of good Y


Here, good X is pork and good y is fish.

a) With income = R40, Themba can consume M/Px amount of good X and M/Py amount of good Y.

Thus, the vertical intercept :

"\frac{40}{5} = 8" units of fish

the horizontal intercept :

"\frac{40}{5} = 5" units of pork







In the diagram, AB is the budget line of Themba, showing Fish on the vertical axis and Pork on the horizontal axis. 


b) Slope of budget line is the ratio of price of x and price of y which is given by marginal rate of exchange. 

Slope of budget line of Themba is Ppork/Pfish

Slope of "AB =\frac{8}{5}"


Answer: Slope of "AB = \frac{8}{5}"


Question 3






i)ATC curve is dependent on AVC and AFC curves. At the beginning the curves fall. The AVC curve will then start to rise as AFC curve continues to fall. The ATC curve will then continue to fall because the falling of AFC curve is greater compared to the rise of AVC curve. The AVC curve will rise as output rises and offset the fall of AFC curve this ATC curve will fall then rise.


ii) The MC curve will fall as output increases and then start to rise at certain output. According to law of variable the marginal cost rises and on reaching maximum, it declines.